2. Calculate the standard deviation and return of portfolio consisting of 60% of Security A and 40% of Security B. Assume cor2.
Answer #1
Ans. 2 For Calculating the Standard Deviation and Return of Portfolio we have to find out return and SD of individual Stock.
Year Return of Security A i.e. RA Return of Security B i.e. RB RA-Mean of RA RA-Mean of RA (RA-Mean of RA)^2 (RA-Mean of RA)^2
2015 20 15 2.8 3.2 7.84 10.24
2016 22 12 4.8 0.2 23.04 0.04
2017 18 11 0.8 -0.8 0.64 0.64
2018 15 10 -2.2 -1.8 4.84 3.24
2019 11 11 -6.2 -0.8 38.44 0.64
TOTAL 86 59 0              -0.00                      74.80                    14.80
For Security A
RA= 86
Mean of RA= ∑RA/n
Where n = Numbers of years i.e. 5
So,
Mean of RA= 86/5
17.2
i.e Expected rate of return from security A = 17.20%
For Security B
RB= 59
Mean of RB= ∑RB/n
59/5
11.8
i.e Expected rate of return from security B = 11.80%
For Security A
Standard Deviation (σA)= (∑(RA-Mean of RA)^2)/n)^1/2
σA= (74.80/5)^1/2
3.87
For Security B
Standard Deviation (σB)= (∑(RB-Mean of RB)^2)/n)^1/2
σB= (14.80/5)^1/2
1.72
As we know that portfolio consisting of 60% of Security A and 40% of Security B
So,
WA=60% i.e. 0.60
WB=40% i.e. 0.40
Mean of RA (i.e. Return from Security A)= 17.2
Mean of RB (i.e. Return from Security B)= 11.8
rAB=0.55
Now
Expected Return of Portfolio = WA*Mean of RA+WB*Mean of RB
= 0.6*17.2+0.4*11.8
15.04 Ans
Expected SD of Portfolio = ( WA*σA)^2+( WB*σB)^2+2*WA*WBAB*rAB
= ((0.60)^2*3.87^2+(0.40)^2*1.72^2+2*0.60*0.40*3.87*1.72*0.55)^1/2
=(7.6223176)^1/2
2.76 Ans

Selecting the portfolio over the individuals stock is better as it wolud reduce the unsysmatic risk.